Write the equation in slope-intercept form of the line that has a slope of -3 and contains the point (1, 1).
Answer by nyc_function(2741)

They give you one point and slope. This is all you need to plug into the point-slope formula.
The point-slope formula is: y - a = m(x - b).
If (a,b) represents your point, then a = 1 and b = 1.
If m represents your slope, then m = -3.
We now plug and chug.
y - 1 = -3(x - 1)
y - 1 = -3x + 3
y = -3x + 3 + 1
y = -3x + 4
Done!